# A hot spots theorem for the mixed eigenvalue problem with small Dirichet region

### v1

@@ 4. Examples
Let $\Omega$ be the square given by the product of intervals $(-1,1)\times(-1,1)$. We first show that if we allow $D$ to have two connected components, then there exist arbitrarily small sets $D$ in $\Omega$ such that $u_{1}^{D}$ has an interior critical point. The second example shows that if we allow $D$ to have four connected components, then there exists arbitrarily small $D$ in $\Omega$ such that $u_{1}^{D}$ has an interior local maximum.
@@
Let $0<\epsilon<\frac{1}{2}$. Let $D_{\epsilon}$ equal the union of the disks $B((\pm\epsilon,0),\epsilon/4)$. See Figure 1. Then the first eigenfunction $u_{1}^{D_{\epsilon}}$ is even about both the $x$- and $y$-axis since it is non-negative. Hence $\partial_{x}u_{1}^{D_{\epsilon}}\equiv 0$ on the $y$-axis and $\partial_{y}u_{1}^{D_{\epsilon}}\equiv 0$ on the $x$-axis. In particular, the origin is a critical point of $u$ for all such $\epsilon$.
@@
Figure 1. The two pairs $(\Omega,D)$ constructed in Example 4.1 each with $\epsilon=0.1$. The dotted lines show the lines of symmetry for each figure. These graphics were created with Desmos (2024).
Now let $D_{\epsilon}$ be the union of the four disks $B((\pm\epsilon,\pm\epsilon),\epsilon/4)$ and $B((\pm\epsilon,\mp\epsilon),\epsilon/4)$. See Figure 1. Since $u$ is non-negative, simple, and even about the lines $y=\pm x$ and the $x$- and $y$-axes, $u$ restricts to a non-negative—and hence, first—mixed eigenfunction of the domain $P=\{(x,y)\mid 0<x<1,0<y<x\}\cap\Omega$ with Dirichlet conditions on the semi-circular arc in $\partial P$. By the argument given for Proposition 2.1 of [JM22], the vertex of $P$ at the origin is a local maximum for the restriction of $u_{1}^{D_{\epsilon}}$ to $P$. This vertex is then also an interior local maximum of $u$.
@@
The case of $n=4$ is a simple computation that we omit (the below computation fails to give the estimate in this case). Suppose that $n\geq 5$. For ease of notation, we will temporarily suppress the subscript $n$ from $d_{n}$. Let $r=r_{n}$ denote the inradius of $P_{n}$ (that is, the supremum of radii of disks contained in $P_{n}$). Suppose that $D$ is contained in the $x$-axis and that $P_{n}$ is contained in the upper half-plane $\{(x,y)\mid y\geq 0\}$. Let $h$ be the length of the orthogonal projection of $P_{n}$ onto the $y$-axis, and let $\ell\geq h$. Let $\varphi(x,y)=\sin\Big{(}\frac{\pi}{2\ell}y\Big{)}$. Note that $\varphi$ is an appropriate test function for the variational characterization (2) of $\lambda_{1}^{D}$.
Let $B_{r}$ be the open disk of radius $r$ centered at $(0,r)$. Because $\varphi$ is increasing in $y$ for $0\leq y\leq h$, we get the lower bound
$\int_{P_{n}}\varphi^{2}\geq\int_{B_{r}}\varphi^{2}.$
We may, without loss of generality, suppose that $P_{n}$ is contained in the rectangle $R=[0,d]\times[0,\ell]$, from which we get the upper bound
$\int_{P_{n}}|\nabla\varphi|^{2}$ $\leq\int_{R}|\nabla\varphi|^{2}=\frac{d\pi^{2}}{8\ell}.$
Combining these estimates gives
(3) $\lambda_{1}^{D}\leq\frac{d\pi^{2}}{8\ell\int_{B_{r}}\sin^{2}\Big{(}\frac{\pi}{2\ell}y\Big{)}}.$
We can scale $P_{n}$ such that $d_{n}=1$. Let $\ell=1$. Then (3) gives the following bound that holds for all $n\geq 5$ (letting $r_{n}$ be the inradius of $P_{n}$):
$\lambda_{1}^{D}\leq\frac{\pi^{2}}{8\int_{B_{r}}\sin^{2}\Big{(}\frac{\pi}{2}y\Big{)}}.$
Since $r_{n}$ is increasing in $n$, we have $B_{r_{n}}\subseteq B_{r_{m}}$ for $n\leq m$, so the right-hand side of this inequality is decreasing in $n$. Hence, it suffices to prove that the right-hand side is less than $j_{0}^{2}$ for $n=5$. Numerical integration performed in Desmos (2024) gives
$\frac{\pi^{2}}{8\int_{B_{r_{5}}}\sin^{2}\Big{(}\frac{\pi}{2}y\Big{)}}<5.4<j_{0}^{2}.$

### v2

@@ 4. Examples
Let $\Omega$ be the square given by the product of intervals $(-1,1)\times(-1,1)$. We first show that if we allow $D$ to have two connected components, then there exist arbitrarily small sets $D$ in $\Omega$ such that $u_{1}^{D}$ has an interior critical point. However, we do not know whether this critical point is a local extremum.
@@
Let $0<\epsilon<\frac{1}{2}$. Let $D_{\epsilon}$ equal the union of the disks $B((\pm\epsilon,0),\epsilon/2)$. See Figure 1. Then the first eigenfunction $u_{1}^{D_{\epsilon}}$ is even about both the $x$- and $y$-axis since it is non-negative. Hence $\partial_{x}u_{1}^{D_{\epsilon}}\equiv 0$ on the $y$-axis and $\partial_{y}u_{1}^{D_{\epsilon}}\equiv 0$ on the $x$-axis. In particular, the origin is a critical point of $u$ for all such $\epsilon$.
@@
Figure 1. The pairs $(\Omega,D)$ constructed in Examples 4.1 and 4.2—each with $\epsilon=0.1$. The dotted lines show the lines of symmetry for each figure. These graphics were created with Desmos (2024).
Example 4.2.
We now show that if $D$ is allowed to have three connected components, then there exists a convex domain with the diameter of $D$ arbitrarily small and such that the first mixed eigenfunction has an interior local extremum. Let $\Omega$ be an equilateral triangle, and let $\epsilon>0$. Number the vertices $v_{1}$, $v_{2}$, $v_{3}$. For each $i=1,2,3$, let $B_{i}$ denote the angle bisector at vertex $v_{i}$. Let $p_{i}$ be the point on $B_{i}$ between $v_{i}$ and the incenter of $\Omega$ such that $p_{i}$ is distance $2\epsilon/\sqrt{3}$ from the incenter. Let $D_{\epsilon}$ equal the union of the closed disks of radius $\epsilon/2$ centered at the $p_{i}$. For $\epsilon$ sufficiently small, $D_{\epsilon}$ is contained in $\Omega$ and has three connected components. See Figure 1. Since $D$ is invariant under the isometry group of $\Omega$ and a first mixed eigenfunction $u_{1}$ does not change signs in $\Omega$, it follows that $u_{1}$ is invariant under the isometry group of $\Omega$. Thus, $u_{1}$ restricts to a first mixed eigenfunction of each fundamental domain $\Omega^{\prime}$ for this group action with Dirichlet conditions on $\Omega^{\prime}\cap D_{\epsilon}$ and Neumann conditions on $\partial\Omega^{\prime}\setminus D_{\epsilon}$. By Proposition 2.1 of [JM22], each acute vertex of $\Omega^{\prime}$ is a local extremum of $u_{1}$. It follows that the incenter of $\Omega$ is a local extremum of $u_{1}$.
We next prove Theorem 1.6, which gives a geometric condition under which a Neumann convex pair $(\Omega,D)$ has a first mixed eigenfunction with no interior critical points.
Proof of Theorem 1.6.
Orient $(\Omega,D)$ in the plane as in the paragraph preceding the theorem statement. Let $R$ denote the rectangle $(0,A/\ell)\times(0,\ell)$. Let $\Omega_{+}=\Omega\cap\{(x,y)\in{\mathbb{R}}^{2}\mid x>0\}$. Then $R$ and $\Omega_{+}$ have the same area, and the area of $R\setminus\Omega_{+}$ has the same area as $\Omega_{+}\setminus R$. Define a test function
$\varphi(x,y)=\begin{cases}0\;\;&\text{if}\;\;x\leq 0\\ \sin\Big{(}\frac{\pi\ell x}{2A}\Big{)}\;\;&\text{if}\;\;0<x<A/\ell\\ 1\;\;&\text{if}\;\;A\leq x.\end{cases}$
Then $\varphi|_{\Omega}\in H^{1}_{D}(\Omega)$. The two estimates
$\int_{\Omega}|\nabla\varphi|^{2}=\int_{\Omega\cap R}|\nabla\varphi|^{2}\leq\int_{R}|\nabla\varphi|^{2}\;\;\text{and}$
$\int_{\Omega}\varphi^{2}=\int_{\Omega_{+}\cap R}\varphi^{2}+\int_{\Omega_{+}\setminus R}\varphi^{2}\geq\int_{\Omega_{+}\cap R}\varphi^{2}+\int_{R\setminus\Omega_{+}}\varphi^{2}=\int_{R}\varphi^{2}$
combine to give
$\lambda_{1}^{D}\leq\frac{\int_{\Omega}|\nabla\varphi|^{2}}{\int_{\Omega}\varphi^{2}}\leq\frac{\int_{R}|\nabla\varphi|^{2}}{\int_{R}\varphi^{2}}=\Big{(}\frac{\pi\ell}{2A}\Big{)}^{2}.$
The sharpness of the estimate and uniqueness of the extremizer follow from the fact that if $\Omega\neq R$, then the inequalities in estimates above become strict. By hypothesis, $d\ell/A\leq 2j_{0}/\pi$, so we have
$\lambda_{1}^{D}\leq\Big{(}\frac{\pi\ell}{2A}\Big{)}^{2}\leq\Big{(}\frac{j_{0}}{d}\Big{)}^{2}.$
The last statement follows from this estimate combined with Proposition 1.5. ∎
Example 4.3.
Let $\Omega$ denote the unit disk, and let $D$ be a subarc of $\partial\Omega$ of length $\alpha\leq\pi$. Then, using the notation of Theorem 1.6, we have
$\frac{d\ell}{A}=\frac{4}{\pi-\alpha/2+\sin(\alpha/2)\cos(\alpha/2)}.$
Numerical computation shows that $d\ell/A\leq 2j_{0}/d$ as long as $\alpha\leq 1.976$. In particular, if $D$ is a subarc of $\partial\Omega$ of length at most $\pi/2$, then $u_{1}^{D}$ has no interior critical points.
Example 4.4.
The definition of Neumann convexity allows Theorem 1.6 to apply to quite wild domains. Let $1<a$, and let $R$ be the rectangle $(0,a-1)\times(0,1)$. Let $R_{-}$ be the closed square $[-1,0]\times[0,1]$. Let $\gamma$ be any simple curve contained in $R_{-}$ joining the points $(0,0)$ and $(0,1)$ that can be locally expressed as the graph of a Lipschitz function. Let $\Omega$ be the region bounded by the edges of $R$ not contained in the $y$-axis and by $\gamma$. See Figure 2. Then
$\frac{d\ell}{A}=\frac{\sqrt{a^{2}+1}}{a-1}\to 1\;\;\text{as}\;\;a\to\infty.$
Thus, for $a$ sufficiently large, $D=\gamma$ defines an eigenvalue problem for which $u_{1}^{D}$ has no critical points.
Figure 2. An example of a Neumann convex pair $(\Omega,D)$ discussed in Example 4.4. The dashed line segments show the square $R_{-}$, and the solid black curve denotes $D=\gamma$. Created with Adobe Illustrator (2024).
@@
Normalize $P_{n}$ so that the distance from the center of $P_{n}$ to each vertex equals $1$. Using the notation of Theorem 1.6, we have
$\frac{d\ell}{A}\leq\frac{4}{A}=\frac{4}{n\sin(\pi/n)\cos(\pi/n)}$
for all $n\geq 4$. The right-hand side of this inequality is decreasing in $n$. One can numerically check that we then have
$\frac{d\ell}{A}\leq\frac{4}{A}<\frac{2j_{0}}{\pi}$
for $n=7$ and therefore for all $n\geq 7$. However, the latter inequality fails to hold for $n\leq 6$, so the cases $n=4,5,6$ require other arguments.
One can check the $n=4$ case directly using the first mixed eigenfunction with $D$ equal to an edge of $P_{4}$ as a test function when $D$ is a subset of an edge.
For $n=5$ with $P_{5}$ normalized as above, we have
$\frac{d\ell}{A}=\frac{2\sin(2\pi/5)(1+\cos(\pi/5))}{5\sin(\pi/5)\cos(\pi/5)}<\frac{2j_{0}}{\pi}$
by numerical computation.
The case $n=6$ is the most involved. We will show directly that $\lambda_{1}^{D}\leq(j_{0}/d)^{2}$ without reference to Theorem 1.6. Rescale, translate, and rotate $P_{6}$ such that $D$ is contained in the $y$-axis, $P_{6}$ has a vertex at the origin, and such that the length $\ell$ of the projection of $P_{n}$ onto the $x$-axis equals $1$. So $P_{6}$ has diameter $2/\sqrt{3}$. Let $R$ be the rectangle $(0,1)\times(0,1/\sqrt{3})$. Let $\tilde{R}$ be the rectangle $(0,1/2)\times(0,1/2\sqrt{3})$. Note that $R\subseteq P_{6}$ (perhaps after reflecting $P_{6}$ about the $y$-axis) and that $P_{6}\setminus R$ has twice the area of $\tilde{R}$. Define a test function $\varphi(x,y)=\sin(\pi x/2)\in H^{1}_{D}(P_{6})$. Then using the monotonicity of $\varphi|_{P_{6}}$ and its derivative in $x$, we have
$\int_{P_{6}}|\nabla\varphi|^{2}\leq\int_{R}|\nabla\varphi|^{2}+2\int_{\tilde{R}}|\nabla\varphi|^{2}=\frac{1}{2\sqrt{3}}\Big{(}\frac{\pi}{2}\Big{)}^{2}\Big{(}\frac{3}{2}+\frac{1}{\pi}\Big{)}\;\;\text{and}$
$\int_{P_{6}}\varphi^{2}\geq\int_{R}\varphi^{2}+2\int_{\tilde{R}}\varphi^{2}=\frac{1}{2\sqrt{3}}\Big{(}\frac{3}{2}-\frac{1}{\pi}\Big{)},$
so we have
$\lambda_{1}^{D}\leq\frac{\int_{P_{6}}|\nabla\varphi|^{2}}{\int_{P_{6}}\varphi^{2}}\leq\Big{(}\frac{\pi}{2}\Big{)}^{2}\cdot\frac{\frac{3}{2}+\frac{1}{\pi}}{\frac{3}{2}-\frac{1}{\pi}}<\Big{(}\frac{j_{0}}{d}\Big{)}^{2},$
where the last inequality holds by direct numerical computation. ∎
@@
Acknowledgements
The author would like to thank Matthew Lowe for illustrating Figure 2. He would also like to thank Mark Ashbaugh for providing several helpful references on the mixed shape optimization problem.